# DyCE: Dynamic Configurable Exiting

### v1

@@ Abstract
Modern deep learning (DL) models necessitate the employment of scaling and compression techniques for effective deployment in resource-constrained environments. Most existing techniques, such as pruning and quantization which primarily aim to generate reduced complexity versions of models through either lowering computational precision or eliminating insignificant elements, are generally static. On the other hand, dynamic compression methods, such as early exits, reduce complexity by recognizing the difficulty of input samples and allocating computation as needed. Dynamic methods, despite their superior flexibility and potential for co-existing with static methods, pose significant challenges in terms of implementation due to any changes in dynamic parts will influence subsequent processes. Moreover, most current dynamic compression designs are monolithic and tightly integrated with base models, thereby complicating the adaptation to novel base models. This paper introduces DyCE, an dynamic configurable early-exit framework that decouples design considerations from each other and from the base model. Utilizing this framework, various types and positions of exits can be organized according to predefined configurations, which can be dynamically switched in real-time to accommodate evolving performance-complexity requirements. We also propose techniques for generating optimized configurations based on any desired trade-off between performance and computational complexity. This empowers future researchers to focus on the improvement of individual exits without latent compromise of overall system performance. The efficacy of this approach is demonstrated through image classification tasks with deep CNNs. DyCE significantly reduces the computational complexity by 23.5% of ResNet152 and 25.9% of ConvNextv2-tiny on ImageNet, with accuracy reductions of less than 0.5%. Furthermore, DyCE offers advantages over existing dynamic methods in terms of real-time configuration and fine-grained performance tuning.
Index Terms:
Network Compression, Model Scaling, Early Exit, Dynamic Network, Deep Learning
I Introduction
@@
Various model compression and optimization strategies, like quantization [2] and pruning [3][4], have been introduced to address this complexity concern. Despite their success to varying degrees, they frequently overlook the impact of the difficulty of input data while optimizing model complexity. In practice, input data can range in difficulty levels, some instances being more challenging than others. Difficult input samples might necessitate deep models for acceptable outcomes, while easier inputs might suffice with smaller (shallower) DL models. However, most existing applications use a static data path for all inputs, thereby wasting computational resources on simpler inputs. As a result, dynamic compression techniques are more suitable for assigning computational resources according to the difficulty of each input sample. Early exit [5, 6] is one of the most promising dynamic techniques due to its straightforward logic and broad flexibility. Typically, several auxiliary exiting networks are connected to the models, aiming to output from intermediate points. The computation ceases at the first exit, where the output satisfies an acceptability criteria.
@@
Figure 1: An overview of DyCE system. (a) is the original deep learning model. (b) is the model when DyCE is applied. The backbone of the original model (shown in red) is divided into segments, and multiple early exits (shown in blue) are attached to the end of each segment. The logic of exiting controllers is illustrated in (c). If the confidence level of an output from the exit point is greater than its associated threshold, the inference will complete immediately. Otherwise, the output of the previous backbone segment will be passed to the next segment. The exit select ($k_{n}$) and corresponding thresholds ($t_{n}$) are from the selected configuration. These configurations are pre-defined to adapt varying performance-complexity targets. Red dots in (d) denote configurations that are available in real-time.
Designing an early exit system necessitates answering three critical questions: (1) How to construct efficient exit networks? (2) Where should the exits be positioned? and (3) When should the system exit? However, solutions to any one of these questions could affect the others, thereby making the design process complex. Furthermore, most existing designs are dedicated to specific base models, often requiring a complete system redesign for the adaptation to any new base models.
@@
In response to these challenges, we introduce DyCE, a Dynamic Configurable Exiting framework. It supports scaling the model in real-time to adapt varying demands. DyCE only requires designers to attach arbitrary exits to an existing base model at arbitrary positions. The search algorithms in the proposed framework will identify the optimal combination of exits and exit conditions for a given performance-complexity target. Thus, questions (2) and (3) are disassociated from question (1), enabling designers to focus on single exit performance without any potential adverse effects on the overall system. Figure 1 provides an overview of the proposed system. The DyCE system can be outlined in three components.
@@
    Early exits (labelled in blue): These are represented by various exit functions attached to the hidden layers of the original model, facilitating an early estimate of the final result. If the output produced at an exit is deemed satisfactory, the inference process concludes at that early layer.
@@
    The exit controller (labelled in green): This is a major difference between DyCE and other early exit systems. It features a configurable exit controller wherein the exit selection and acceptance criteria are controlled by pre-configured configurations generated via search algorithms to meet any performance-complexity targets.
@@
    DyCE Framework: We introduce DyCE, a configurable early-exit-based framework, which can be applied and retrofitted to any existing DL models so as to scale and compress the model dynamically.
@@
    Real-time Adaptability: The proposed approach enables run-time selection of complexity-accuracy tradeoff points to adapt to varying practical demands. Furthermore, it is compatible with a wide range of deep learning models and other compression methods.
@@
    Decoupled Design: DyCE decouples the design considerations for creating an early-exit system, simplifying the design process. This design separation allows for focused improvements on individual exits while ensuring systematic performance by DyCE.
@@
    Search Algorithms: The proposal of two search algorithms for generating configurations suitable for DyCE. These configurations encompass optimized plans detailing which exiting functions to employ and their associated thresholds, catering to different trade-off targets.
@@
The rest of this paper is structured as follows: Section II reviews the related compression and scaling methods. Section III discusses the inference process proposed in DyCE. Section IV explores considerations in the design of exit networks. Section V introduces the search algorithm proposed for generating configurations. Finally, the experimental results are discussed in Section VI.
II Related Work
TABLE I: Feature comparison between the proposed method and related work
[b] Category Method Work No Extra Network Loosely Coupled1 No Backbone Training Required Post-training Configurable Real-time Configurable Fine-grained Tuning Dynamic Early Exit DyCE × BranchyNet[5] × × × ×2 ×2 ×2 MSDNet[8] MSNet[9] × × × ×2 ×2 ×2 Chen et al.[10] EPNet[11] × × × × × × Layer Skipping SkipNet[12] × × × × × × Channel Skipping RNP[13] × × × × × × Static Architecture Scaling[14][15] × × × × × Quantization[16] × × Pruning[3] × ×
@@ II Related Work
    The method is not majorly relying on specific model architecture or task.
@@
    Support is possible if our proposed method is applied.
Model compression and scaling can be achieved using static or dynamic methods. Static methods typically apply reductions to model structure and weights, while dynamic methods adapt to each individual input sample [17]. We have reviewed and categorized mainstream model compression and scaling techniques, summarizing the primary distinctions in Table I.
II-A Static Methods
@@
Architecture Scaling, an integral part of modern model architectures, allows for the scaling of architectures into different versions [14][15] by modifying the number of repeating components. However, this scaling is coarse-grained, and each version of the model must be trained from scratch.
Quantization [16] is a prevalent compression technique that uses fewer bits to represent network weights or activations. Pruning [3][4] is another common approach that removes redundant weights from a trained neural network. Both these methods can be applied during pre-training or post-training phases, thereby allowing for compressing models to different scales without the need for retraining.
II-B Dynamic Methods
@@
Dynamic compression methods allocate computational resources based on the complexity of each input sample. The Early Exit strategy is a dynamic compression technique that enables “easy” samples to exit at shallow layers, avoiding the execution of deeper network segments. This concept was initially introduced by BranchyNet [5] with confidence-based exits, where computation terminates if the confidence of a classifier exceeds a threshold. Subsequent studies such as FastBERT [18] and PersEPhonEE [19] have extended this confidence-based method across various applications. In these methods, the exit propensity is controlled by external thresholds, thereby enabling post-training or real-time configuration. Alternatively, some research uses policy networks rather than thresholds to control exits [10][11]. These policy networks are trained for specific compression targets, offering improved performance, albeit with less flexibility after training completion.
Other work, such as MSDNet [8] and MSNet [9], achieves high performance by enabling communication among exits at different resolutions. Compared to standard fixed network models, these approaches offer dynamic compression but necessitate task-specific designs. i.e. these approaches cannot be used to retrofit scalability into an existing base model network.
Apart from early exits, which skip all subsequent layers, Layer Skipping [12] is a method that skips some intermediate layers while retaining the remaining model segments. There are also strategies involving Channel Skipping [13], which omits only certain parts of a layer during runtime. Designing and training models that can adapt to skipped intermediate parts present significant challenges, making these methods less prevalent than the early exit approach.
III Inference with Early Exits
An early-exit system can be implemented in various ways. In this section, we outline the inference process employed by DyCE. The runtime component of DyCE is a rapid inference system based on early exits, which dynamically reduces overall computation during model inference. This system facilitates exits at multiple layers on the base model in accordance with a pre-defined configuration. When inference computation reaches at an exit point, the process will terminate if the model exhibits confidence in the preliminary prediction. Hence, more computations are allocated to inputs that cause the exits to have less confidence regarding early-exit decisions, i.e. difficult inputs, while easy inputs are mostly exited at early stages. Additionally, this system can switch configurations at runtime to achieve dynamic scaling.
III-A Runtime Architecture
@@
To illustrate the run-time architecture of DyCE, we start with a pre-trained network (i.e. the base model), e.g., ResNet. Further, the model is divided into different segments, and exit networks are attached to each of them. These exits are trained and grouped into various configurations to obtain the required trade-off. These ideas are now formally defined in the following sub-sections.
@@ III-A2 Exit
where $\tilde{\boldsymbol{y}}_{n,k}$ has a format compatible [^1] Note that we define that $\tilde{\boldsymbol{y}}_{n,k}$ is compatible with the original network’s output but not necessarily in the exact same format. It could, for example, be shorter, providing classification results for a subset of classes or perhaps some clustering of classes. with the original network’s output. This is illustrated in Fig. 2. For ease of description, we consider the pre-trained small network at the end of the original backbone network as being another exit, denoted $f_{N,1}$. Any input sample that does not exit early will eventually pass through this final exit. Each of the exit function is designed and trained offline independently of any other and are fixed at run-time.
III-A3 Confidence and Threshold
The exit confidence is an estimate of the correctness of the result. For classification problems, we take this confidence to be the maximum value of the predicted class probability, i.e., $max(\tilde{\boldsymbol{y}}_{n,k})$. In our system, we apply a threshold, $t$, to this confidence value in order to make an early-exit decision. A high threshold limits an exit such that only highly confident samples exit early, and a lower threshold will allow more samples to exit early. Accordingly, the threshold, $t$, can be adjusted to obtain different complexity/accuracy trade-offs as will be discussed in Section V.
III-A4 Exit Group and Run-time Algorithm
Our system uses pre-defined exit configuration groups, each denoted as $G$, that enumerates which exit functions and what threshold values are to be used at the output of each segment. Specifically we define an exit configuration, $G=\{\boldsymbol{k},\boldsymbol{t}\}$, comprising a length $N$ ordered list of function indices, $\boldsymbol{k}$, and thresholds, $\boldsymbol{t}$, such that the exit to be used after the $n^{\text{th}}$ segment is $f_{n,k_{n}}$, i.e. $\boldsymbol{\tilde{y}}_{n}\triangleq\boldsymbol{\tilde{y}}_{n,k_{n}}$ and the corresponding confidence is $\max(\tilde{\boldsymbol{y}}_{n})$. This confidence is then compared against the threshold $t_{n}$ in order to make an early-exit decision. If the confidence is not less than the threshold, i.e. $\max(\tilde{\boldsymbol{y}}_{n})\geq\boldsymbol{t}_{n}$, the computation is terminated (early exited) and the $\tilde{\boldsymbol{y}}_{n}$ vector is returned. Note that we set $\boldsymbol{k}_{N}=1$ and $\boldsymbol{t}_{N}=0$ forcing the last exit to always be $f_{N,1}$, and forcing all samples exit eventually. This run-time algorithm is summarized in the algorithm 1.

### v3

@@ Abstract
Conventional deep learning (DL) model compression methods affect all input samples equally. However, as samples vary in difficulty, a dynamic model that adapts computation based on sample complexity offers a novel perspective for compression and scaling. Despite this potential, existing dynamic techniques are typically monolithic and have model-specific implementations, limiting their generalizability as broad compression and scaling methods. Additionally, most deployed DL systems are fixed, and unable to adjust once deployed. This paper introduces DyCE, a dynamically configurable system that can adjust the performance-complexity trade-off of a DL model at runtime without needing re-initialisation or re-deployment. DyCE achieves this by adding exit networks to intermediate layers, thus allowing early termination if results are acceptable. DyCE also decouples the design of exit networks from the base model itself, enabling its easy adaptation to new base models. We also propose methods for generating optimized configurations and determining exit network types and positions for desired trade-offs. By enabling simple configuration switching, DyCE provides fine-grained performance tuning in real-time. We demonstrate the effectiveness of DyCE through image classification tasks using deep convolutional neural networks (CNNs). DyCE significantly reduces computational complexity by 26.2% for ResNet152, 26.6% for ConvNextv2tiny and 32.0% for DaViTbase on ImageNet validation set, with accuracy reductions of less than 0.5%.
keywords:
[…]
{highlights}
Effectively compress the computational complexity of deep learning models.
Enable early exiting on any existing deep learning models by attaching tiny exits.
Dynamically change the exiting configuration to achieve arbitrary scaling and adapt the varying demand in real-time.
Generating the best exiting configuration of a multi-exit system for various trade-off preferences.
1 Introduction
@@
Various model compression and optimization strategies, like quantization [2] and pruning [3][4], have been introduced to address this complexity concern. Despite their success to varying degrees, they overlook the impact of the difficulty of input data while optimizing model complexity. In practice, input data can range in difficulty levels, some instances being more challenging than others. Difficult input samples might necessitate deep models for acceptable outcomes, while easier inputs might suffice with smaller (shallower) DL models. However, most existing applications use a static data path for all inputs, thereby wasting computational resources on simpler inputs. Routing input samples to various sub-models based on their difficulty level is better suited for allocating computational resources efficiently. Early exit [5, 6] is one of the most promising dynamic techniques due to its simplicity and flexibility. Typically, several auxiliary exiting networks are connected to the base models, generating outputs from intermediate points in the base model. The computation ceases at the first exit, where the output satisfies an acceptability criterion. However, most early exit based dynamic models are monolithic designs. i.e. Their components, such as backbone models, exit branches, and exit policies, are tightly coupled. This coupling makes it challenging to apply these designs to other or future models. Additionally, these designs often sacrifice some performance to incorporate dynamic features, limiting their effectiveness as compression methods.
@@
Figure 1: An overview of DyCE system. (a) is the original deep learning model. (b) is the model when DyCE is applied. The backbone of the original model (shown in red) is divided into segments, and multiple early exits (shown in blue) are attached to the end of each segment. The logic of exiting controllers is illustrated in (c). If the confidence level of an output from the exit point is greater than its associated threshold, the inference will complete immediately. Otherwise, the output of the previous backbone segment will be passed to the next segment. The exit select ($k_{n}$) and corresponding thresholds ($t_{n}$) are from the selected configuration. These configurations are pre-defined to adapt varying performance-complexity targets.
Furthermore, most DL systems, particularly those with static models, encounter a common issue: once designed and deployed in a system, their compression ratio and performance are fixed. Adjustments require redeploying another model. In real-world applications, performance requirements can vary based on environmental or operational conditions. This issue persists for many dynamic models [7, 8, 9, 10, 11], as their routing strategies are trained. Therefore, they are not configurable post-training or in real-time. Although some other dynamic models [5, 12, 13, 14, 15] allow configuring their routing strategies externally, they typically use a single threshold across all exits to modify the performance-complexity trade-off. Given that different exits are relatively independent and have suitable working conditions individually, this uniform approach is rigid and inefficient for systems with multiple exits. A more flexible routing strategy is needed to achieve optimal real-time scaling.
This work introduces DyCE, a Dynamically Configurable Exiting framework. DyCE is an early-exit DL framework that can dynamically configure the performance-complexity trade-off point of a DL model during run-time, without re-initialising or re-deploying a updated model on an inference hardware. DyCE enhances existing base models by adding arbitrary exit points at any position, eliminating the necessity for re-training or redesigning the original model. Unlike other efficient designs that create new dynamic models, DyCE can transform static models into dynamic ones. This capability allows the latest advanced deep learning models, which are typically static, to benefit from dynamic data path. The flexibility of DyCE in both the system building and running phases makes it a versatile method for model compression and scaling. Designing an early exit system necessitates answering three challenging questions: (1) How to construct efficient exit networks? (2) Where should the exits be positioned? and (3) When should the system exit? However, solutions to any one of these questions could affect the others, thereby making the design process complex. Furthermore, most existing designs are dedicated to specific base models, often requiring a complete system redesign for the adaptation to any new base models.
@@
The search algorithms in the proposed framework will identify the optimal combination of exits and exit conditions for a given performance-complexity target. Thus, questions (2) and (3) are disassociated from question (1), enabling designers to focus on single exit performance without any potential adverse effects on the overall system. Fig. 1 provides an overview of the proposed system. The DyCE system can be outlined in three components.
@@
    Early exit functions (labelled in blue): These are represented by various functions attached to the hidden layers of the original model, facilitating an early estimate of the final result. If the output produced at an exit is deemed satisfactory, the inference process concludes at that early layer.
@@
    The exit controller (labelled in green): This is a major difference between DyCE and other early exit systems. The $n^{\text{th}}$ exit is controlled by $k_{n}$ and $t_{n}$, means that exit function $f_{n,k_{n}}$ is used with the threshold $t_{n}$. It features a configurable exit controller wherein the exit selection and acceptance criteria are controlled by pre-configured configurations generated via search algorithms to meet any performance-complexity targets.
In the proposed system, we transform a static model (Fig. 1 (a)) into a dynamic model (Fig. 1 (b)) by adding additional networks at numerous potential positions while maintaining the integrity of the original model. As shown in (Fig. 1 (c)), the exiting behavior is controlled by the exit function index $k_{n}$ and the threshold $t_{n}$ for each position. Modifying these variables changes the system’s performance. The specific values for each $k_{n}$ and $t_{n}$ are searched configurations. The search algorithm discussed in Sec. 5 provides configurations for different preferences in fine-grain. During the runtime, the system loads exits and apply the thresholds according to $k_{n}$ and $t_{n}$ of a specific configuration. The dynamic scalability is achieved by switching between different configurations. Fig. 2 is an example of possible configurations (marked as red dots) that DyCE creates for a base model. The density of options can be further increased by using a smaller searching step while creating configurations.
Figure 2: An example of DyCE with ResNet50. All red dots selectable performance-complexity operating points in real-time.
@@
    DyCE Framework: The introduced DyCE framework can be applied and dynamically scale and compress existing DL models.
@@
    Computational Saving: DyCE can reduce the computation by 26.2% for ResNet152, 26.6% for ConvNextv2tiny and 32.0% for DaViTbase on ImageNet, with accuracy reductions of less than 0.5%.
@@
    Real-time Scaling: Supporting the selection of complexity-accuracy points to adapt to real-time demands flexibly.
@@
    Decoupled Design: DyCE decouples the design considerations for creating an early-exit system, simplifying the developing process.
  5. 5.
    Search Algorithms: Search algorithms are also proposed to generate configurations for different tradeoff preferences.
@@
The rest of this paper is structured as follows: Sec. 2 reviews the related compression and scaling methods. Sec. 3 discusses the inference process proposed in DyCE. Sec. 4 explores considerations in the design of exit networks. Sec. 5 introduces the search algorithm proposed for generating configurations. Finally, the experimental results are discussed in Sec. 6.
2 Related Work
Table 1: Feature comparison between the proposed method and other compression or scaling methods
[b] Category Method Work Scale Up/Down Loosely Coupled1 No Required Backbone Training Post-training Configurable1 Real-time Configurable1 Fine-grained Tuning1 Utilizing dynamic models Confidence-based Early Exit DyCE Down BranchyNet[5] MSDNet[12] MSNet[13] DVT[14] CF-ViT[15] Down × × -2 -2 -2 Network-controlled Early Exit Chen et al.[7] EPNet[8] Down × × × × × Layer Skipping SkipNet[9] Down × × × × × Channel Skipping RNP[10] Down × × × × × Sparse MoE[11] V-MoE[17]Mobile V-MoE[18] Both3 × × × × × General Architecture Scaling[19][20] Both × × × × Knowledge distillation[21] Down × × × Pruning[3] Down 4 × Quantization[22] Down 4 × ×
@@ 2 Related Work
    Loosely coupled refers the method is not majorly relying on specific model architecture or task. Post-training configurable means the compression rate can be modified without further training. Real-time configurable means the compression rate can be modified during the inference runtime. Fine-grained Tuning means the method can achieve arbitrary compression instead of selecting from limited choices.
@@
    They meet the prerequisites for implementing these features. Support is possible if our proposed method is applied.
  • 3
    Sparse MoE is more often known as an efficient up scaling method, but created efficient models can be considered as down scaled models.
  • 4
    Backbone training after pruning or quantization is also common, but not necessary.
Model compression and scaling refer to reducing model costs with acceptable performance loss or improving model performance efficiently with increased costs. Most compression methods produce static, immutable models with a fixed compression rate once deployed. Dynamic models, however, can adapt to individual input samples [23]. The dynamic nature of these models offers a new dimension for compression and scaling, providing difficulty-aware compression and configurable compression ratios through a dynamic computation graph. We reviewed and categorized model compression and scaling techniques, including both static methods and those utilizing dynamic models, summarizing the primary distinctions in Tab. 1.
2.1 Compression & Scaling for General Models
@@
Architecture Scaling is an integral part of modern model architectures, allows for the scaling of architectures into different versions [19][20] by modifying the number of repeating components. However, this scaling is coarse-grained, and each version of the model must be trained from scratch.
Quantization[22] is a prevalent compression technique that uses fewer bits to represent network weights or activations. Quantization cannot reduce the number of operations but can achieve some speedups with specific hardware supports because quantized variables theoretically only require simpler computing units[24].
Pruning[3][4] is another common approach that removes redundant weights from a trained neural network. However, pruning usually introduces sparsity which may be inefficient for hardware execution.
Knowledge distillation[21] is a method to create a smaller model with the teaching of a large model. Unlike Quantization or Pruning which can be applied to both pre-training and post-training phases, Knowledge distillation necessitates a massive training to teach the small model with large model’s predictions.
2.2 Compression & Scaling Methods Utilizing Dynamic Models
@@
Dynamic models make it possible to allocate computational resources based on the complexity of each input sample. The Early Exit strategy that enables “easy” samples to exit at shallow layers, avoiding the execution of deeper network segments. This concept was initially introduced by BranchyNet[5] with confidence-based exits, where computation terminates if the confidence of a classifier exceeds a threshold. Subsequent studies such as FastBERT [25] and PersEPhonEE[26] have extended this confidence-based method across various applications. In these methods, the exit propensity is controlled by external thresholds, thereby enabling post-training tuning. Alternatively, some research uses policy networks rather than thresholds to control exits [7][8]. These policy networks are trained for specific compression targets, offering improved performance, albeit with less flexibility after training completion. Other work, such as MSDNet[12], MSNet[13], DVT[14] and CF-ViT[15], achieves high performance by enabling feature reusing. Compared to efficient static models, these approaches offer efficiency by differentiating sample difficulties but necessitate task-specific designs. i.e. these approaches cannot be used to retrofit scalability into an existing base model network. Apart from early exits, which skip all subsequent layers, Layer Skipping[9] is an idea that skips some intermediate layers while retaining the remaining model segments. There are also strategies involving Channel Skipping[10], which omits only certain parts of a layer during runtime. Designing and training models that can adapt to skipped intermediate parts present significant challenges, making these methods less prevalent than the early exit approach. Sparse Mixture of Experts (MoE)[11, 17, 18] is another wide-sued model scaling technique. It employs multiple experts (sub-models in parallel) but adaptively activate only some of them during inference. Typically, experts in MoE implementations possess equal complexity,making Sparse MoE cannot address sample difficulties by varying the amount of computation. Nevertheless, this approach significantly expands model capacity by substantially increasing the parameter count while maintaining consistent computational costs.
3 Inference with Early Exits
We start illustrating our system with the inference of a model with multiple early exits. Before the inference, we attach multiple extra networks to a base model as exit points and generate various configurations to coordinate all exits. We describe how to achieve these in Sec. 4 and Sec. 5.
DyCE reduces overall computation by dynamically routing samples based on their difficulty. When inference reaches an exit point, the process terminates if the model’s confidence exceeds a predefined threshold. Otherwise, computation continues to the next backbone segment. This approach allocates more resources to difficult inputs, through trigger later exits, while easy inputs are processed through early exits. Additionally, DyCE can switch configurations at runtime to adjust the exit ratio at a macro level. If more accuracy is requested, more samples will be routed to later exits with more computation. Conversely, for low computational cost is more important, samples will exit earlier during inference.
3.1 Runtime Architecture
@@
To illustrate the run-time architecture of DyCE, we divide a pre-trained network (i.e. the base model), e.g., ResNet into different segments, and exit networks are attached to each of them. These exits are trained and grouped into various configurations to obtain the required trade-off. These ideas are now formally defined in the following sub-sections.
@@ 3.1.2 Exit
where $\tilde{\boldsymbol{y}}_{n,k}$ has a format compatible [^1] Note that we define that $\tilde{\boldsymbol{y}}_{n,k}$ is compatible with the original network’s output but not necessarily in the exact same format. It could, for example, be shorter, providing classification results for a subset of classes or perhaps some clustering of classes. with the original network’s output. As illustrated in Fig. 3, we provide multiple exit candidates at the same position. The configuration searching algorithm (discussed in Sec.5) will identify the best one for a given compression preference, and only that one will be loaded during the inference. In our experiments, we use basic multilayer perceptron (MLP)[27] with different layers and neurons, but they can be any designs align with the definition above. For ease of description, we consider the pre-trained small network at the end of the original backbone network as being another exit, denoted $f_{N,1}$. Any input sample that does not exit early will eventually pass through this final exit. Each of the exit function is designed and trained offline independently of any other and are fixed at run-time.
3.1.3 Confidence and Threshold
The exit confidence is an estimate of the correctness of the result. For classification problems, we take this confidence to be the maximum value of the predicted class probability, i.e., $max(\tilde{\boldsymbol{y}}_{n,k})$. In our system, we apply a threshold, $t$, to this confidence value in order to make an early-exit decision. A high threshold limits an exit such that only highly confident samples exit early, and a lower threshold will allow more samples to exit early. Accordingly, the threshold, $t$, can be adjusted to obtain different complexity/accuracy trade-offs as will be discussed in Sec. 5.
3.1.4 Exit Group and Run-time Algorithm
Our system uses pre-defined exit configuration groups, each denoted as $G$, that enumerates which exit functions and what threshold values are to be used at the output of each segment. Specifically we define an exit configuration, $G=\{\boldsymbol{k},\boldsymbol{t}\}$, comprising a length $N$ ordered list of function indices, $\boldsymbol{k}$, and thresholds, $\boldsymbol{t}$, such that the exit to be used after the $n^{\text{th}}$ segment is $f_{n,k_{n}}$, i.e. $\boldsymbol{\tilde{y}}_{n}\triangleq\boldsymbol{\tilde{y}}_{n,k_{n}}$ and the corresponding confidence is $\max(\tilde{\boldsymbol{y}}_{n})$. This confidence is then compared against the threshold $t_{n}$ in order to make an early-exit decision. If the confidence is not less than the threshold, i.e. $\max(\tilde{\boldsymbol{y}}_{n})\geq t_{n}$, the computation is terminated (early exited) and the $\tilde{\boldsymbol{y}}_{n}$ vector is returned. Note that we set $k_{N}=1$ and $t_{N}=0$ forcing the last exit to always be $f_{N,1}$, and forcing all samples exit eventually. This run-time algorithm is summarized in the Algorithm 1.